# State-space models are accurate and efficient neural operators for dynamical systems

### v1

@@ 4.3 Out-of-Distribution Operator Learning Tasks
Regarding model structure, LSTM and GRU are one-layer models with 64 hidden dimensions. DeepONet’s trunk and branch nets are all five-layer networks with 24 hidden dimensions. FNO is a three-layer model with 32 hidden dimensions and 8 modes. Regarding LNO, we opt for two versions. The first is the original model from Cao et al. [4], whose model size is smaller (LNO-small), and the other is a larger one-layer LNO with 16 hidden dimensions and 8 modes (LNO-large). We opt for a four-layer Transformer with 24 hidden dimensions, 96 feedforward dimensions, and 4 heads. For all three Oformers with vanilla, Galerkin, or Fourier attention, we opt for three-layer Oformers with 24 hidden dimensions, 24 feedforward dimensions, and 4 heads. GNOT has three layers, 16 hidden dimensions and 16 feedforward dimensions, 4 heads, and 2 experts. Regarding Mamba, we opt for two blocks with skip connection and layer normalization, with 16 model dimensions and 16 intermediate dimensions.
Model Params Memory Time Lorenz-5 Lorenz-10 Duffing-0 Duffing-.5 Pend-0 Pend-.5
LSTM 33537 1965MB 16min 6.381E-3 5.163E-1 1.000E+0 1.283E-1 7.589E-1 1.352E-1
GRU 25217 1909MB 16min 1.276E-2 4.789E-1 8.089E-1 1.282E-1 7.065E-1 1.159E-1
DeepONet 103176 1385MB 5min 1.216E-1 7.933E-1 9.136E-1 5.887E-1 1.145E+0 5.773E-1
FNO 27873 1693MB 8min 2.662E-2 3.769E-1 5.895E-1 1.938E-1 6.835E-1 1.544E-1
LNO (Small) 1309 1885MiB 1min 1.748E-2 1.420E-1 2.598E-1 1.125E-1 2.278E-1 1.624E-1
LNO (Large) 6737 1987MB 7min 1.982E-1 8.006E-1 9.988E-1 9.805E-1 1.200E+0 9.309E-1
Transformer 29041 20849MB 232min 1.434E-2 2.877E-1 1.241E+0 1.092E-1 7.754E-1 8.497E-2
Oformer-V 33505 9607MB 162min 4.264E-2 2.634E-1 5.110E-1 1.996E-1 8.701E-1 2.888E-1
Oformer-G 33361 1949MB 37min 3.131E-2 8.388E-1 1.049E+0 2.948E-1 9.137E-1 2.317E-1
Oformer-F 33361 1949MB 37min 2.704E-2 7.216E-1 9.896E-1 2.566E-1 9.993E-1 2.926E-1
GNOT 33479 2333MB 78min 3.259E-2 8.738E-1 1.122E+0 2.856E-1 9.392E-1 3.285E-1
Mamba (Ours) 19233 1635MB 10min 5.956E-3 2.754E-1 4.445E-1 8.957E-2 8.978E-1 8.837E-2
@@ 4.4 Long-Time Integration
We aim to learn the operator $f(t)\mapsto x(t)$. We choose the damping $c\in\{0.1,0.3,0.5\}$ and $g/l=1$. We also wish to test different models’ stability under various damping factors $c$. We manipulate the input and output sequence lengths in 2048, 4096, 8192, 16384, and 32768, corresponding to terminal times in 20.48, 40.96, 81,92, 163.84, and 327.68. The original setting in Cao et al. [4] only considers terminal time up to 20.48, while we consider a much longer time. We test the out-of-distribution (OOD) generalization problem following Cao et al. [4], i.e., the training and testing functions/data will follow different distributions. Specifically, we sample 200 training samples from $f_{train}(t)=A\sin(5t)$, where the $A\in\{0.05,0.1,0.15,\cdots,9.95,10\}$. We sample 50 validation samples and 130 testing samples from $f_{test}(t)=A\exp(-0.001t)\sin(5t)$, where $A\in\{0.14,0.19,\cdots,9.04,9.09\}$. The temporal discretization step size is 0.01. The implementation details, such as model structure, training procedure, and test metrics, are the same as those in Section 4.3.
@@
Table 5 presents different models’ scalability and speed, while Table 6 presents the performance comparison results. Regarding scalability presented in Table 5, we observe a linear growth in running time and memory cost for RNNs, linear Transformers (including Oformer-G, Oformer-F, and GNOT), and Mamba. On the other hand, Transformer and Oformer-V, with a conventional quadratic self-attention scheme, exhibit fast growth in memory and running time and cannot scale up to sequences longer than 8192, i.e., their memory costs exceed the limit of the used NVIDIA A100 GPU with 81292MB memory. This demonstrates Mamba’s linear complexity while keeping the model expressive and parallelizable. Furthermore, the neural operators’ training time and memory cost are more stable with input sequence lengths since they disregard temporal information. However, this also leads to their inferior performance, as explained below. Regarding performance comparisons presented in Table 6, we report N.A. for models that require running longer than 10 hours (600 minutes) since it exceeds a reasonable training time or models going out-of-memory. Mamba performs best in most cases, especially when dealing with long-term problems. Mamba is only outperformed by other models in the pendulum problem with damping $c=0.1$ and sequence lengths 2048 and 4096, while Mamba is the best in all other settings. Even in test cases where it is outperformed by other models, Mamba’s performance is still close to the best one. Mamba’s errors grow more mildly than RNNs with the sequence length/terminal time, possibly because Mamba is better at capturing long-term dependencies than RNNs, which are known to forget. The errors of Linear Transformers grow with sequence length, reflecting a tradeoff in Transformers’ computational complexity and model expressiveness. More specifically, the Transformer’s costly self-attention scheme is motivated by capturing long sequences. Despite their efficiency, linear Transformers approximate the original quadratic self-attention trade-off computational complexity with the capability to capture long-term dependencies, leading to their fast error growth concerning sequence length. On the other hand, Mamba combines the advantages of both being efficient with linear cost and being expressive enough to capture long-time dependencies. Regarding neural operators, we also observed a fast growth in their errors with the sequence length because they are general neural operators failing to capture the temporal information intrinsic in dynamical systems. For instance, DeepONet’s model parameters grow linearly with sequence length since it directly inputs the more extended input sequence. FNO and LNO transform the input function using trainable kernel integration operators on a bounded domain. With a longer time, FNO and LNO compress the original temporal domain $[0,T]$ to a bounded domain, say $[0,1]$. This domain transformation will cause the input function to contain more high-frequency signals, e.g., the long-frequency function $\sin(x)$ can be transformed to a high-frequency one $\sin(Tx)$ when $T$ is large. Consequently, this makes the training and generalization more challenging for FNO and LNO, leading to their inferior performances under long-time integration.
@@
Sequence Length 2048 4096 8192 16384 32768
Terminal Time $T$ 20.48 40.96 81.92 163.84 327.68
GRU 2.479E-2 2.640E-2 1.199E-1 1.410E-1 1.654E-1
LSTM 3.273E-2 3.919E-2 1.049E-1 9.521E-2 9.319E-2
DeepONet 7.124E-1 7.527E-1 7.837E-1 8.308E-1 8.623E-1
FNO 2.022E-2 3.518E-2 4.273E-2 1.299E-1 3.483E-1
LNO 1.051E-1 1.435E-1 8.787E-1 9.722E-1 9.827E-1
Transformer 4.488E-2 N.A. N.A. N.A. N.A.
Oformer-V 1.740E-2 N.A. N.A. N.A. N.A.
Oformer-G 2.590E-2 6.169E-1 7.211E-1 Diverge Diverge
Oformer-F 2.368E-2 5.817E-1 6.980E-1 Diverge Diverge
GNOT 2.870E-2 4.378E-2 7.693E-2 N.A. N.A.
Mamba 1.785E-2 3.181E-2 3.803E-2 3.415E-2 3.461E-2
@@
Sequence Length 2048 4096 8192 16384 32768
Terminal Time $T$ 20.48 40.96 81.92 163.84 327.68
GRU 5.249E-3 7.939E-3 1.009E-2 1.496E-2 1.309E-2
LSTM 7.350E-3 7.854E-3 7.168E-3 6.867E-3 5.743E-2
DeepONet 5.994E-1 6.785E-1 7.781E-1 8.416E-1 8.954E-1
FNO 7.346E-3 1.001E-2 1.237E-2 4.259E-2 6.783E-2
LNO 8.079E-2 2.535E-1 7.661E-1 9.687E-1 9.904E-1
Transformer 1.773E-2 N.A. N.A. N.A. N.A.
Oformer-V 9.266E-3 N.A. N.A. N.A. N.A.
Oformer-G 1.527E-2 7.951E-2 1.438E-1 Diverge Diverge
Oformer-F 1.326E-2 7.819E-2 1.195E-1 Diverge Diverge
GNOT 1.066E-2 1.995E-2 3.889E-2 N.A. N.A.
Mamba 4.725E-3 4.465E-3 5.327E-3 4.543E-3 4.992E-3
@@
Sequence Length 2048 4096 8192 16384 32768
Terminal Time $T$ 20.48 40.96 81.92 163.84 327.68
GRU 3.325E-3 4.434E-3 4.053E-3 6.893E-3 3.371E-2
LSTM 3.222E-3 3.399E-3 3.203E-3 7.658E-3 4.803E-3
DeepONet 6.375E-1 7.199E-1 8.296E-1 8.859E-1 9.286E-1
FNO 4.633E-3 8.257E-3 1.264E-2 8.723E-2 1.476E-1
LNO 1.423E-1 6.970E-2 5.413E-1 9.605E-1 9.915E-1
Transformer 1.278E-2 N.A. N.A. N.A. N.A.
Oformer-V 8.938E-3 N.A. N.A. N.A. N.A.
Oformer-G 1.210E-2 5.695E-2 1.468E-1 Diverge Diverge
Oformer-F 1.086E-2 5.301E-2 1.584E-1 Diverge Diverge
GNOT 9.174E-3 1.483E-2 2.197E-2 N.A. N.A.
Mamba 2.131E-3 2.490E-3 2.533E-3 3.862E-3 2.030E-3
@@
Figure 5: Visualization of Mamba’s prediction on the pendulum problem in Cao et al. [4] for long-time integration corresponding to Section 4.4 in this paper. (a) $T\in[0,20.48]$. (b) $T\in[20.48,40.96]$. (c) $T\in[40.96,61.44]$. (d) $T\in[266.24,286.72]$. (e) $T\in[286.72,307.2]$. (f) $T\in[307.2,327.68]$. The relative $L_{2}$ error with respect to time is plotted in Figure 6. The performance and scalability comparisons between different models are presented in Tables 5 and 6.
Figure 6: Visualization of Mamba’s relative $L_{2}$ prediction error concerning time in the long-time pendulum problem taken from Cao et al. [4] corresponding to Section 4.4 of this paper. (a) $c=0.1$. (b) $c=0.3$. (c) $c=0.5$. The performance and scalability comparisons between different models are presented in Tables 5 and 6. We visualize Mamba’s prediction in Figure 5.
We plot Mamba’s prediction versus the ground truth on test data 0 in the Pendulum problem with $c=0.5$ and the most extended sequence, 32768, in Figure 5. The ODE dynamics is complicated at the beginning. However, as time evolves and with the damping effect, the solution gradually converges to a sinusoidal-like function. Since Mamba achieves a relative $L_{2}$ error of about 1e-3 in this case, the ground truth and its prediction are almost indistinguishable, further validating Mamba’s outstanding accuracy. We plot the relative $L_{2}$ error with respect to time for Pendulum problems with $c=0.1,0.3,0.5$ and the most extended sequence, 32768, for LSTM, GRU, and Mamba, in Figure 6, which further validates the above observation. Specifically, the errors are larger at the beginning due to the relatively more complex dynamics there, as shown in the figure. At later times, the dynamics gradually converge to a sinusoidal-like function, while the errors drop and maintain at a steady level.

### v2

@@ 4.3 Out-of-Distribution Operator Learning Tasks
We select model structures based on the reported validation dataset. We set a parameter count budget of approximately 100,000. The model search space is as follows. For time series models, including GRU, LSTM, Mamba, Transformer, Oformer, and GNOT, the search space is model depth in $\{1,2,3,4\}$, hidden dimensions in $\{8,12,16,24,32,48,64,96,128\}$. We set all Transformer-related models’ head numbers as four and the number of experts in GNOT as two. We set DeepONet’s trunk and branch nets as MLP and search the MLP depth in $\{2,3,4,5\}$, and a corresponding hidden size satisfies the parameter budget. We search FNO’s depth in $\{2,3,4\}$, hidden dimensions in $\{8,16,32\}$ and number of modes in $\{4,8,16\}$. We search LNO’s depth in $\{1,2,3\}$, hidden dimensions in $\{4,8,16\}$ and number of modes in $\{4,8,16\}$. After validation, we choose the following model structures that perform the best among all the six validation sets on average. LSTM and GRU are one-layer models with 64 hidden dimensions. DeepONet’s trunk and branch nets are all five-layer networks with 24 hidden dimensions. FNO is a three-layer model with 32 hidden dimensions and 8 modes. Regarding LNO, we opt for two versions. The first is the original model from Cao et al. [4], whose model size is smaller (LNO-small), and the other is a larger one-layer LNO with 16 hidden dimensions and 8 modes (LNO-large). We opt for a four-layer Transformer with 24 hidden dimensions, 96 feedforward dimensions, and 4 heads. For all three Oformers with vanilla, Galerkin, or Fourier attention, we opt for three-layer Oformers with 24 hidden dimensions, 24 feedforward dimensions, and 4 heads. GNOT has three layers, 16 hidden dimensions and 16 feedforward dimensions, 4 heads, and 2 experts. Regarding Mamba, we opt for two blocks with skip connection and layer normalization, with 16 model dimensions and 16 intermediate dimensions.
Model Params Memory Time Lorenz-5 Lorenz-10 Duffing-0
LSTM 33537 1965MiB 16min 6.381E-3$\pm$7.821E-4 5.163E-1$\pm$1.029E-1 1.000E+0$\pm$6.831E-2
GRU 25217 1909MiB 16min 1.276E-2$\pm$4.674E-3 4.789E-1$\pm$1.313E-1 8.089E-1$\pm$6.008E-2
DeepONet 103176 1385MiB 5min 1.216E-1$\pm$2.392E-2 7.933E-1$\pm$2.554E-1 9.136E-1$\pm$8.952E-2
FNO 27873 1693MiB 8min 2.662E-2$\pm$1.930E-3 3.769E-1$\pm$1.002E-1 5.895E-1$\pm$6.538E-2
LNO (Small) 1309 1885MiB 1min 1.748E-2$\pm$3.571E-3 1.420E-1$\pm$2.796E-2 2.598E-1$\pm$3.298E-2
LNO (Large) 6737 1987MiB 7min 1.982E-1$\pm$5.609E-2 8.006E-1$\pm$2.635E-1 9.988E-1$\pm$9.895E-2
Transformer 29041 20849MiB 232min 1.434E-2$\pm$4.190E-3 2.877E-1$\pm$1.133E-1 1.241E+0$\pm$1.780E-1
Oformer-V 33505 9607MiB 162min 4.264E-2$\pm$2.508E-3 2.634E-1$\pm$3.451E-2 5.110E-1$\pm$2.189E-1
Oformer-G 33361 1949MiB 37min 3.131E-2$\pm$2.412E-3 8.388E-1$\pm$8.587E-2 1.049E+0$\pm$6.729E-2
Oformer-F 33361 1949MiB 37min 2.704E-2$\pm$1.722E-3 7.216E-1$\pm$9.066E-2 9.896E-1$\pm$1.153E-1
GNOT 33479 2333MiB 78min 3.259E-2$\pm$1.735E-3 8.738E-1$\pm$9.707E-3 1.122E+0$\pm$4.003E-2
Mamba (Ours) 19233 1635MiB 10min 5.956E-3$\pm$1.493E-3 2.754E-1$\pm$2.045E-1 4.445E-1$\pm$2.297E-1
Model Params Memory Time Duffing-.5 Pend-0 Pend-.5
LSTM 33537 1965MiB 16min 1.283E-1$\pm$7.865E-3 7.589E-1$\pm$1.036E-1 1.352E-1$\pm$3.432E-3
GRU 25217 1909MiB 16min 1.282E-1$\pm$5.121E-3 7.065E-1$\pm$2.152E-1 1.159E-1$\pm$5.147E-3
DeepONet 103176 1385MiB 5min 5.887E-1$\pm$1.203E-1 1.145E+0$\pm$3.175E-2 5.773E-1$\pm$9.989E-2
FNO 27873 1693MiB 8min 1.938E-1$\pm$1.727E-2 6.835E-1$\pm$1.921E-1 1.544E-1$\pm$2.423E-2
LNO (Small) 1309 1885MiB 1min 1.125E-1$\pm$2.098E-2 2.278E-1$\pm$4.556E-2 1.624E-1$\pm$2.074E-2
LNO (Large) 6737 1987MiB 7min 9.805E-1$\pm$2.546E-1 1.200E+0$\pm$2.873E-1 9.309E-1$\pm$2.357E-1
Transformer 29041 20849MiB 232min 1.092E-1$\pm$1.777E-2 7.754E-1$\pm$2.416E-1 8.497E-2$\pm$2.770E-2
Oformer-V 33505 9607MiB 162min 1.996E-1$\pm$1.725E-2 8.701E-1$\pm$3.055E-1 2.888E-1$\pm$2.850E-2
Oformer-G 33361 1949MiB 37min 2.948E-1$\pm$2.902E-2 9.137E-1$\pm$2.995E-2 2.317E-1$\pm$3.961E-2
Oformer-F 33361 1949MiB 37min 2.566E-1$\pm$4.633E-2 9.993E-1$\pm$4.672E-2 2.926E-1$\pm$1.445E-2
GNOT 33479 2333MiB 78min 2.856E-1$\pm$3.886E-2 9.392E-1$\pm$1.594E-2 3.285E-1$\pm$1.992E-2
Mamba (Ours) 19233 1635MiB 10min 8.957E-2$\pm$1.366E-2 8.978E-1$\pm$3.487E-2 8.837E-2$\pm$8.949E-3
@@ 4.4 Long-Time Integration
We aim to learn the operator $f(t)\mapsto x(t)$. We choose the damping $c\in\{0.1,0.3,0.5\}$ and $g/l=1$. We also wish to test different models’ stability under various damping factors $c$. We manipulate the input and output sequence lengths in 2048, 4096, 8192, 16384, and 32768, corresponding to terminal times in 20.48, 40.96, 81,92, 163.84, and 327.68. The original setting in Cao et al. [4] only considers terminal time up to 20.48, while we consider a much longer time. We test the out-of-distribution (OOD) generalization problem following Cao et al. [4], i.e., the training and testing functions/data will follow different distributions. Specifically, we sample 200 training samples from $f_{train}(t)=A\sin(5t)$, where the $A\in\{0.05,0.1,0.15,\cdots,9.95,10\}$. We sample 50 validation samples and 130 testing samples from $f_{test}(t)=A\exp(-0.001t)\sin(5t)$, where $A\in\{0.14,0.19,\cdots,9.04,9.09\}$. The temporal discretization step size is 0.01. The implementation details, such as model structure, training procedure, and test metrics, are the same as those in Section 4.3 due to the similar data generation setup.
@@
Table 6 presents different models’ scalability and speed, while Table 7 presents the performance comparison results. Regarding scalability presented in Table 6, we observe a linear growth in running time and memory cost for RNNs, linear Transformers (including Oformer-G, Oformer-F, and GNOT), and Mamba. On the other hand, Transformer and Oformer-V, with a conventional quadratic self-attention scheme, exhibit fast growth in memory and running time and cannot scale up to sequences longer than 8192, i.e., their memory costs exceed the limit of the used NVIDIA A100 GPU with 80GB memory. This demonstrates Mamba’s linear complexity while keeping the model expressive and parallelizable. Furthermore, the neural operators’ training time and memory cost are more stable with input sequence lengths since they disregard temporal information. However, this also leads to their inferior performance, as explained below. Regarding performance comparisons presented in Table 7, we report N.A. for models that require running longer than 10 hours (600 minutes) since it exceeds a reasonable training time or models going out-of-memory. Mamba performs best in most cases, especially when dealing with long-term problems. Mamba is only outperformed by other models in the pendulum problem with damping $c=0.1$ and sequence lengths 2048 and 4096, while Mamba is the best in all other settings. Even in test cases where it is outperformed by other models, Mamba’s performance is still close to the best one. Mamba’s errors grow more mildly than RNNs with the sequence length/terminal time, possibly because Mamba is better at capturing long-term dependencies than RNNs, which are known to forget. The errors of Linear Transformers grow with sequence length, reflecting a tradeoff in Transformers’ computational complexity and model expressiveness. More specifically, the Transformer’s costly self-attention scheme is motivated by capturing long sequences. Despite their efficiency, linear Transformers approximate the original quadratic self-attention trade-off computational complexity with the capability to capture long-term dependencies, leading to their fast error growth concerning sequence length. On the other hand, Mamba combines the advantages of both being efficient with linear cost and being expressive enough to capture long-time dependencies. Regarding neural operators, we also observed a fast growth in their errors with the sequence length because they are general neural operators failing to capture the temporal information intrinsic in dynamical systems. For instance, DeepONet’s model parameters grow linearly with sequence length since it directly inputs the more extended input sequence. FNO and LNO transform the input function using trainable kernel integration operators on a bounded domain. With a longer time, FNO and LNO compress the original temporal domain $[0,T]$ to a bounded domain, say $[0,1]$. This domain transformation will cause the input function to contain more high-frequency signals, e.g., the long-frequency function $\sin(x)$ can be transformed to a high-frequency one $\sin(Tx)$ when $T$ is large. Consequently, this makes the training and generalization more challenging for FNO and LNO, leading to their inferior performances under long-time integration.
@@
Seq. Length 2048 4096 8192 16384 32768
Time $T$ 20.48 40.96 81.92 163.84 327.68
GRU 2.479E-2$\pm$2.561E-3 2.640E-2$\pm$3.008E-3 1.199E-1$\pm$1.854E-2 1.410E-1$\pm$2.691E-2 1.654E-1$\pm$3.178E-2
LSTM 3.273E-2$\pm$4.558E-3 3.919E-2$\pm$5.719E-3 1.049E-1$\pm$1.089E-2 9.521E-2$\pm$1.227E-2 9.319E-2$\pm$1.325E-2
DeepONet 7.124E-1$\pm$6.178E-2 7.527E-1$\pm$9.381E-2 7.837E-1$\pm$1.399E-1 8.308E-1$\pm$1.403E-1 8.623E-1$\pm$1.875E-1
FNO 2.022E-2$\pm$8.840E-4 3.518E-2$\pm$8.963E-4 4.273E-2$\pm$1.465E-3 1.299E-1$\pm$1.855E-3 3.483E-1$\pm$3.590E-3
LNO 1.051E-1$\pm$3.547E-2 1.435E-1$\pm$4.097E-2 8.787E-1$\pm$8.266E-2 9.722E-1$\pm$7.454E-2 9.827E-1$\pm$1.267E-1
Transformer 4.488E-2$\pm$3.284E-3 N.A. N.A. N.A. N.A.
Oformer-V 1.740E-2$\pm$9.853E-4 N.A. N.A. N.A. N.A.
Oformer-G 2.590E-2$\pm$1.720E-3 6.169E-1$\pm$5.541E-2 7.211E-1$\pm$5.965E-2 Diverge Diverge
Oformer-F 2.368E-2$\pm$1.998E-3 5.817E-1$\pm$6.874E-2 6.980E-1$\pm$9.506E-2 Diverge Diverge
GNOT 2.870E-2$\pm$5.266E-3 4.378E-2$\pm$8.089E-3 7.693E-2$\pm$1.286E-2 N.A. N.A.
Mamba 1.785E-2$\pm$6.106E-3 3.181E-2$\pm$8.263E-3 3.803E-2$\pm$8.615E-3 3.415E-2$\pm$7.128E-3 3.461E-2$\pm$6.988E-3
@@
Seq. Length 2048 4096 8192 16384 32768
Time $T$ 20.48 40.96 81.92 163.84 327.68
GRU 5.249E-3$\pm$9.880E-4 7.939E-3$\pm$1.562E-3 1.009E-2$\pm$1.778E-3 1.496E-2$\pm$3.741E-3 1.309E-2$\pm$3.268E-3
LSTM 7.350E-3$\pm$2.064E-3 7.854E-3$\pm$2.775E-3 7.168E-3$\pm$1.957E-3 6.867E-3$\pm$1.276E-3 5.743E-2$\pm$8.745E-3
DeepONet 5.994E-1$\pm$7.845E-2 6.785E-1$\pm$8.233E-2 7.781E-1$\pm$8.547E-2 8.416E-1$\pm$1.067E-1 8.954E-1$\pm$1.854E-1
FNO 7.346E-3$\pm$8.772E-4 1.001E-2$\pm$6.945E-4 1.237E-2$\pm$1.451E-3 4.259E-2$\pm$3.337E-3 6.783E-2$\pm$8.976E-3
LNO 8.079E-2$\pm$8.815E-3 2.535E-1$\pm$1.745E-2 7.661E-1$\pm$5.676E-2 9.687E-1$\pm$9.763E-2 9.904E-1$\pm$1.050E-1
Transformer 1.773E-2$\pm$2.945E-3 N.A. N.A. N.A. N.A.
Oformer-V 9.266E-3$\pm$1.089E-3 N.A. N.A. N.A. N.A.
Oformer-G 1.527E-2$\pm$2.062E-3 7.951E-2$\pm$4.712E-3 1.438E-1$\pm$1.046E-2 Diverge Diverge
Oformer-F 1.326E-2$\pm$2.556E-3 7.819E-2$\pm$3.561E-3 1.195E-1$\pm$7.892E-3 Diverge Diverge
GNOT 1.066E-2$\pm$2.385E-3 1.995E-2$\pm$3.686E-3 3.889E-2$\pm$5.371E-3 N.A. N.A.
Mamba 4.725E-3$\pm$1.288E-3 4.465E-3$\pm$1.293E-3 5.327E-3$\pm$1.760E-3 4.543E-3$\pm$1.885E-3 4.992E-3$\pm$1.925E-3
@@
Seq. Length 2048 4096 8192 16384 32768
Time $T$ 20.48 40.96 81.92 163.84 327.68
GRU 3.325E-3$\pm$5.966E-4 4.434E-3$\pm$4.850E-4 4.053E-3$\pm$4.988E-4 6.893E-3$\pm$5.022E-4 3.371E-2$\pm$4.741E-3
LSTM 3.222E-3$\pm$6.288E-4 3.399E-3$\pm$6.796E-4 3.203E-3$\pm$4.475E-4 7.658E-3$\pm$6.296E-4 4.803E-3$\pm$2.953E-4
DeepONet 6.375E-1$\pm$6.974E-2 7.199E-1$\pm$8.205E-2 8.296E-1$\pm$7.236E-2 8.859E-1$\pm$7.784E-2 9.286E-1$\pm$8.006E-2
FNO 4.633E-3$\pm$9.870E-4 8.257E-3$\pm$1.455E-3 1.264E-2$\pm$3.809E-3 8.723E-2$\pm$1.066E-2 1.476E-1$\pm$3.964E-2
LNO 1.423E-1$\pm$3.871E-2 6.970E-2$\pm$1.055E-2 5.413E-1$\pm$8.674E-2 9.605E-1$\pm$1.048E-1 9.915E-1$\pm$1.474E-1
Transformer 1.278E-2$\pm$6.871E-4 N.A. N.A. N.A. N.A.
Oformer-V 8.938E-3$\pm$1.065E-3 N.A. N.A. N.A. N.A.
Oformer-G 1.210E-2$\pm$4.508E-4 5.695E-2$\pm$8.773E-4 1.468E-1$\pm$1.986E-2 Diverge Diverge
Oformer-F 1.086E-2$\pm$5.267E-4 5.301E-2$\pm$9.880E-3 1.584E-1$\pm$2.352E-2 Diverge Diverge
GNOT 9.174E-3$\pm$2.054E-3 1.483E-2$\pm$3.833E-3 2.197E-2$\pm$3.568E-4 N.A. N.A.
Mamba 2.131E-3$\pm$7.819E-4 2.490E-3$\pm$8.020E-4 2.533E-3$\pm$8.114E-4 3.862E-3$\pm$1.185E-3 2.030E-3$\pm$5.640E-4
@@
Figure 7: Visualization of Mamba’s prediction on the pendulum problem in Cao et al. [4] for long-time integration corresponding to Section 4.4 in this paper. (a) $T\in[0,20.48]$. (b) $T\in[20.48,40.96]$. (c) $T\in[40.96,61.44]$. (d) $T\in[266.24,286.72]$. (e) $T\in[286.72,307.2]$. (f) $T\in[307.2,327.68]$. The relative $L_{2}$ error with respect to time is plotted in Figure 8. The performance and scalability comparisons between different models are presented in Tables 6 and 7.
Figure 8: Visualization of Mamba’s relative $L_{2}$ prediction error concerning time in the long-time pendulum problem taken from Cao et al. [4] corresponding to Section 4.4 of this paper. (a) $c=0.1$. (b) $c=0.3$. (c) $c=0.5$. The performance and scalability comparisons between different models are presented in Tables 6 and 7. We visualize Mamba’s prediction in Figure 7.
We plot Mamba’s prediction versus the ground truth on test data 0 in the Pendulum problem with $c=0.5$ and the most extended sequence, 32768, in Figure 7. The ODE dynamics are complicated at the beginning. However, as time evolves and with the damping effect, the solution gradually converges to a sinusoidal-like function. Since Mamba achieves a relative $L_{2}$ error of about 1e-3 in this case, the ground truth and its prediction are almost indistinguishable, further validating Mamba’s outstanding accuracy. We plot the relative $L_{2}$ error with respect to time for Pendulum problems with $c=0.1,0.3,0.5$ and the most extended sequence, 32768, for LSTM, GRU, and Mamba, in Figure 8, which further validates the above observation. Specifically, the errors are larger at the beginning due to the relatively more complex dynamics there, as shown in the figure. At later times, the dynamics gradually converge to a sinusoidal-like function while the errors drop and maintain a steady level.